OpenStudy (anonymous):

What is the area of a circle with a radius of 10 feet? (Use 3.14 for ρ.)

OpenStudy (anonymous):

A=pi r^2 A=(3.14)(10^2) A=3.14(100) A=314 square feet

OpenStudy (anonymous):

u want the answer in which unit